FEL Upgrade Project Weekly Brief – October 14-18, 2002




We have had a full week of installation and commissioning activities on the FEL upgrade. Highlights include: (1) successful powering of the injector dipole magnets; (2) completion and sign-off of the operating procedure to allow us to power the SRF modules and execution of this procedure began today; (3) completion and installation of the photogun light box assembly; (4) receipt of the ultrahigh vacuum chambers for the optical cavity mirror assemblies; and (5) successful re-coating of the photogun ball electrode with the field emission suppression coating.

We remind our readers that Jefferson Lab will be hosting the 2002 Industrial Physics Forum for the American Institute of Physics at Williamsburg on Oct. 27-29th. This meeting will highlight speakers on Monday Oct. 28th from recent users of the FEL during the 2000-2001 user run, and on Tuesday, Oct. 29th a session on "Science and Security" will include the President’s Science Advisor, John Marburger, and the Director of Science and Technology from the Office of Naval Research, Jane Alexander. The full program for the meeting is given at www.aip.org/ipf.


There is full calendar for the next month on top of the installation and commissioning activities. Preparations had to be interleaved for next week’s DOE/SURA review of "Institutional Management", final preparation for the Oct. 27-29 Industrial Physics Forum noted above; preparation of papers for the AVS Symposium, Denton Symposium on Applications of Accelerators and the DEPS Symposium (all in November); and initial preparations for our next semiannual review on Nov. 20-21.

F.Dylla attended Virginia Research and Technology Advisory Commission held on Monday, Oct. 14th in Mclean,VA on behalf of the Director. We learned this week that our state funds that support FEL operations will be cut 15% effective on Oct. 1 because of the state’s budget short fall. 2/3 of the state budget lines were cut at this level. On the positive side, the state administration is willing to put the FEL Building addition back into a bond issue for consideration by the General Assembly in January.

Michelle Shinn attended a portion of the International Congress on Applications of Lasers and Electro-Optics

WBS 4 (Injector):

The leak in a weld on the gun chamber was repaired and the chamber is under vacuum on an ion pump with screens, anode plate, NEG pumps and the instrument flange installed. The rear stalk drive bellows was welded and is complete. The solenoid with correctors nested inside was mounted on the gun’s downstream vacuum tube. The rear external corona dome for the top of the ceramic stack assembly was received. Assembly of the light box was completed and it was installed onto its stand in the vault.

The first of two implant runs on the ball cathode was successfully completed on Wednesday. The wafer side has been implanted first, resulting in a very uniform pink colored coating. The stainless steel support mount as well as the ball cathode are free of the dark contamination previously observed with the aluminum support mount. The implant run on the back side was successfully completed late this afternoon. The ball will slowly cool in a dry nitrogen environment over the weekend. It will be removed

on Monday to resume final assembly of the electrode structure.

Gun HVPS - The current meters were calibrated for the Remote Control units. The resistor modification to achieve 600 kV is in progress. The bellows connecting the HVPS Tank to the Gun Tank has been drilled, dogs ears welded to the flanges, returned, and cleaned. It will be installed next week. All of the parts are in house for the combined Conditioning & Running resistors mechanism. Final assembly is in progress

WBS 6 (RF):

Zone 2 - Ready for SRF checkout.

Zone 4 - Additional trimming is needed for the waveguides. The Machine Shop has promised them for early next week. The PSS certification that was scheduled for Friday noon was not successful. Waveguide pressure would not hold those parts that were installed this week. Leak checking is in progress. RadCon had the wrong detector and its cable was not routed properly. A Lock & Tag was left on the CPS by someone out of town. Everything will be corrected by next Tuesday and PSS will be tried again.

Quarter HVPS - Some progress with the EPICS software, but it is still a week or so away from being operational. The HVPS and klystron cart for unit 4 are still not complete. People were used to install waveguides in zone 4.

WBS 8 (Instrumentation):

This week was the time for dotting the "I"s and crossing the "T"s. The final cleanup for many of the installed systems is underway. Installation and termination of the corrector magnets in the 2F and 3F regions is almost complete. Installation and checkout of the corrector magnets in the 0F region is accelerating to meet scheduling goals. Testing of the 5 kW Injection Dipole PS for the GV magnets proved successful.

BLMs (Beam Loss Monitors) - High level software is making good progress. The diagnostics screens are functional and the application screens are actively being checked out. As the prototype boards are coming to the end of the development stage, a list of assembly and installation notes is being compiled to be included with the boards documentation package. Current EPICS screen shots are here: -




Drive Laser Controls - The variable attenuator motor control system will be operated by a 'new' 4-channel controller card. This card is on order. The changes to the MPS shutter interface chassis are still ongoing. Additional modifications to the Acromag have been made to allow the DAC output to be bi-polar for the EO cells. The optics group will be running the laser today and certifying the work that has been done. Misc. - An ECO on the beam viewer camera/lamp interface board is in progress. The changes will address the certain PCB mounting and interconnect issues that have complicated their reliability. Additionally, several other PCB CAD package check prints are ready to be reviewed. These include the VME interface daughter board (F0117), the 8-channel fiber driver (F0118), the 16-channel isolation/relay I/O module


Fast Shutdown Down interlock and cabling for RF in zone 4 was installed this week. Supporting software for FSD is complete and checkout is progress. The 15 pin fanout boards for the 64-bit Interface Module have been received from the vendor and being populated in house.

Modifications to the Dump Skid junction box have been completed to support the new Conductivity Analyzers from Solu Comp. Rebuild of the chassis is progressing.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• The re-ground core pieces for the remaining 6 dipoles met specification. They report excellent results. They are gluing the Purcell Gap sheet pieces down to the core plates.

• Two of four GWs are on the comparison test stand for their first formal measurements. After much effort by lots of folks, a glitch was found by the Accelerator electronics support folks. Repair allowed the comparison probes to work. The second feature of the stand, the loop coil, which detects field flatness, is still giving problems but a plan is in place to commission it. We are looking foreword to successful measurements next week.

Injector Dipoles (DU/DV)

• The first GU was measured. The current is about 6 A. lower than the GV to get the proper core field. The shunt will take care of that. Success! The core field is flat to 2 parts in 10000 over the good field region. A second style of magnet made it. The first grid to find the field integral flatness was taken. That data is not looked at by press time.

• The first string of 3 GVs was wired up and powered in the injector line.

Arc 180 Degree Dipoles (GY)

• Wang NMR finished winding the last GY coil.

• Wang stopped gluing down the Purcell Gap Shim pieces. The person in charge is taking a vacation back to his homeland in China. We are about to work this issue.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Wang vacuum pressure impregnated the 4th GX coil to complete the set of coils for the first arc. They are assembling two GQs and two GXs and plan to ship the four magnets for the First Arc next week.


Trim Quad (QT)

• Contrary to last week’s report, the second batch of 6 GQs was not qualified on the stand. Two were qualified this week and two were found to be miswired. We will correct them and continue testing.

Sextupole (SF)

• The contract was awarded to Master Machine to disassemble the eight Sextupoles from the IR Demo (SCs) and add pieces to and double them up and make four "Sextupole Light" magnets (Designated "SGs" as a substitute for the SFs.)

Octupole (OT)

• We are working on the specification to go with the drawing.

Beam Line and Vacuum

• We hit a small snag in our contract with Master Machine on the chambers. We are getting a discoloration around the weld zone on the inside even though the chamber is well purged. This led to the discovery that the chambers still have an oil contamination on the inside. The procedure we taught them of washing down the surfaces with solvents before "clean welding for vacuum" has degraded in practice. Master Machine masterfully responded by steam cleaning the tacked up 180 degree chamber that was about to be welded over the weekend. They will steam clean (a much more effective cleaning method) all parts before welding in the future.

• Other pipes in the injection line continued to be installed by the Vacuum Group.

• Y Chamber – It is being leak tested.

• Chicane chambers started up fabrication again. One is tacked up for final welding and another is almost through part machining. Two more are in preparation.

• Girder and vacuum assembly continued.

• The girder drawings for around the wiggler region (Last for the IR machine) are ready for sign off.

• The last of the stands (except for those left out for dipole magnet installation) were aligned.

• The magnet measurement crew installed many of the correctors on girders.

WBS 11 (Optics):

We received the optical cavity vacuum vessels for the IR Upgrade. They are being uncrated, and will be worked on, one at a time, in our clean enclosure in User Lab 2. There is a part that needs to be welded to the upper assembly, this is due in today (10/18). We anticipate the top hats will be ready for reintegration as early as ~ 10 days from now. We have spent several days looking at the optical cavity hardware internals to determine if they were cleaned to our specifications. Some parts are, but others aren't, so it appears we'll have to disassemble the internals for cleaning. The Survey and Alignment group has established the mounting coordinates for the optical cavities and their associated pedestals and girders for the alignment diagnostics. The anchors are drilled, and the pedestals installed and aligned. The OCMMS hardware is ready for installation and one beamline assembly was installed today. Hardware to establish a clean environment around the optical cavity vessels was received and will be installed once alignment activities in the area are complete. The trolley hoists were received - once installed they will be used to lift the vacuum vessel "top hats" on and off the lower section. This will make assembly and testing much easier. The optical transport turning mirrors (turning cassettes 1, 3, & 4) are being checked and detailed. The holder for the 3" outcouplers was signed and being procured. This is also true for the insertable mirror. All parts for the 0F region THz diagnostic were received. They were test assembled then sent to be UHV cleaned. They will be installed in the X-chamber at the earliest opportunity.


A test procedure (TSOP) was written and approved to allow us to do SRF commissioning on off shifts while we continue installation and non-powered check-out activities during the first shift. We thank the installation crew for preparing new portable lead shields for this activity. SRF commissioning should commence next week.

Details of the week’s commissioning activities are given in the WBS 4,6,8 and 9 write-ups above.

Last week's test of the drive laser EPICS controls turned up some issues in implementation. These are being resolved.